Suppose that x and y are int variables.
The statement x=++y+5; is equivalent to the statement(s)_________

a. x=y+5;
b. x=y+6;
c. x=y+1;
d.None of these
Answer #1

Answer: B

  • Because in statement y is pre-incrementing so y will incremented by 1. Then adding 5. so x = (y+1)+5=y+6.
  • So the answer is B. ie x=y+6.